# A Survey on Tabular Data Generation: Utility, Alignment, Fidelity, Privacy, and Beyond

### v1

@@ 2 Synthesising Tabular Data
In what follows, we group existing deep generative modelling methods for synthesising tabular data by the requirement they primarily address.
@@
3 Utility
Determining whether synthetic data can replace real training data in downstream tasks has been the most common evaluation approach (e.g., [????]). The main reason for this is that synthetic data are often generated to augment or create datasets for real-world applications, where good predictive performance is crucial, e.g., predicting diabetes diagnoses from ICU data [?]. Therefore, utility still remains a key requirement for synthetic data and benchmark for comparing synthesisers.
3.1 Evaluation Protocol and Metrics
The most common protocol used to evaluate the utility of generative models is the Train on Synthetic, Test on Real (TSTR) Protocol (see, e.g., [?]). The procedure is to train a suite of ML models (e.g., a support vector machine, XGBoost) on the synthetic data, and then test their performance on a real test set, reporting the performance using different metrics on the ground of the task defined by the dataset of interest: for classification (resp., regression) datasets, standard metrics include accuracy and F1 score (resp., root mean square error and explained variance).
3.2 Methods
@@
GAN-based models. Many of the works surveyed here are based on the original Generative Adversarial Network (GAN) [?], which relies on two neural networks: (i) a generator, which given a noise vector returns a synthetic datapoint and hence learns the real data distribution, and (ii) a discriminator, which given a data point (either real or synthetic) classifies it as either real or synthetic and hence learns to distinguish between synthetic and real data. The GAN jointly trains these two models in an adversarial framework, where the generator’s objective is to produce samples that confuse the discriminator. See the first schema from the left in Figure 1 for a simple abstraction of GAN-based models. CWGAN: Building on the seminal Wasserstein GAN (WGAN) [?], which used the Wasserstein distance as a loss function for a GAN model, ? [?] proposed CWGAN as an oversampling framework for class balancing. CWGAN generates synthetic samples for underrepresented classes, augmenting the training set. Results indicate that oversampling is beneficial primarily for strongly non-linear datasets. OCT-GAN: proposed by ? [?], incorporates neural ordinary differential equations (neural ODEs) in their GAN through the introduction of a new layer that allows for the extraction of a sequence of hidden vector representations given an input sample, i.e., the hidden evolution trajectory of the sample. The trajectory is used to help the discriminator decide whether a sample is real or synthetic. While improving utility, the usage of ODEs makes OCT-GAN slower than most GAN-based models.
GAN and BN-based models. GANBLR: ? (?) construct the generator and discriminator as Bayesian networks (BNs), which are able to encode known feature interactions. The background knowledge inclusion leads to better utility than standard GANs.
GAN and flow-networks-based models. TabGFN: ? (?) (i) model conditional distributions between the features using flow networks [?], where the features are ordered and sequentially generated, making it possible to trace the generation process and construct a sampling process trajectory; and (ii) adopt the WGAN-GP [?] critic with gradient penalty as a reward function to learn these relationships and the order of features. A key limitation is that TabGFN cannot handle continuous features, which can affect performance on high-dimensional continuous data having large variance.
Diffusion-based models. TabDDPM: Motivated by the their success in computer vision, ? ? introduced TabDDPM: a diffusion-based model for synthesising tabular data. Popular for its high utility results, TabDDPM is able to model both discrete and continuous features, but separately, by using multinomial [?] and Gaussian diffusion models [?], respectively. For a simple abstraction on how a diffusion model can be used to generate data, see the central schema in Figure 1. STaSy: Proposed by ? [?], STaSy is based on a score-based generative model (SGM), which was shown to be a competitive alternative to diffusion-based models in [?]. Contrarily to TabDDPM, StaSy estimates the score function (gradient of log-probability) instead of explicitly modelling the forward and reverse Markov processes. As the loss is difficult to train, STaSy uses a self-paced learning strategy, which starts with a subset of training data—yielding a stable, low loss—and gradually expands to the full dataset. Evaluated on 15 real-world datasets against 7 baselines (mostly GAN-based models), STaSy achieves high utility. Due to the forward and backward passes, both StaSy and TabDDPM suffer from slow sampling times. Moreover, the base SGM model is known to be unstable in high-dimensional settings. CoDi: Motivated by the difficulty of modelling discrete features, ? [?] proposed CoDi, a framework comprising co-evolving diffusion models that separately handle continuous and discrete features as in TabDDPM [?], with the difference that the two models condition on each other during training. At each training step, the discrete (resp., continuous) diffusion model takes the perturbed sample from the continuous (resp., discrete) one as input, and both models are conditioned on both samples from the previous step during the denoising process. Slower than most GAN-based models (excluding OCT-GAN), CoDi is faster than STaSy and TabDDPM, while also obtaining higher utility on mixed types datasets.
LLM-based models. GReaT [?] represents an answer to the abundance of works that convert tabular data into numerical representations thus losing the contextual connections between the features, which very often carry semantic meaning. Indeed, GReaT follows these steps (also illustrated in Figure 1): first, the tabular data undergoes a textual encoding process, converting it into text. Next, a feature order permutation step is applied. The resulting sentences are then used to fine-tune the large language models (LLMs) [?]. At sampling time, either a single feature name or arbitrary feature-value pairs are given as input to the LLM, which then completes them (note that this allows for arbitrary conditioning). Using this approach, GReaT outperforms most of the baselines w.r.t. utility.
Energy-based models. TabPFGen [?] is a generative model that uses a pretrained network as an energy-based model for data augmentation and class balancing. Although its capabilities are limited to low-dimensional inputs (it was tested on datasets with maximum 10 features), TabPFGen demonstrates good utility performance in the data augmentation task by simply using a pretrained model compared to specialised GAN-, VAE-, and diffusion-based models.
4 Alignment
Brought to the attention of the research community more recently [?] than the other requirements, background knowledge alignment is an important condition for synthetic data when evaluating its realism.
4.1 Evaluation Protocol and Metrics
In [?], three metrics have been proposed to evaluate alignment: (i) the Constraint Violation Rate (CVR), which computes the percentage of samples that do not satisfy at least one of the constraints in the available set of constraints, (ii) Constraint Violation Coverage (CVC), computing the percentage of constraints that have been violated at least once by the sample set, and (iii) the sample-wise Constraint Violation Coverage (sCVC) which determines the average percentage of samples violating each of the constraints. Out of these three metrics, CVR has also been used in [?], albeit called feasibility rate there.
4.2 Methods
GAN and AE-based models. ITS-GAN: proposed for tabular data augmentation ITS-GAN [?] is an early method advocating for the use of background knowledge. The knowledge it can express is of two types: (i) rules stating that the value of one feature uniquely determines the value of another (e.g., the feature Position determines the feature Salary), and (ii) rules stating that a specific value for a set of features determines the specific value for another (e.g., if Position = “CEO” then Salary = “2M”). To encode the first types of rules, they train on the real data an autoencoder (AE) for each rule to output the value of the dependent features given the values of the independent ones. Then, at training time, the GAN generator is trained to minimise the difference between its output and one of the AEs, while also being penalised if it violates any rule of the second type. Additionally, the discriminator receives as input the difference between the AE’s outputs and the synthetic samples, thus being able to use this information to assess the sample.
GAN and BN-based models. C3-TGAN: Similarly to GANBLR, C3-TGAN [?] uses Bayesian networks to capture background knowledge. However, while GANBLR models only discrete features, C3-TGAN can handle both discrete and continuous features by following the CTGAN [?] approach. To incorporate explicit attribute correlations and property constraints from background knowledge, it represents constraints as control vectors (similar to the conditional vectors in CTGAN) and uses them to guide training, ensuring better alignment.
VAE and GNN-based models. GOGGLE: [?] is a different approach which integrates knowledge about pairwise feature dependencies by encoding them in a graph where each feature is a node and an edge exists if a relation between the two features is known. New relations can be learnt using a message passing mechanism from graph neural networks (GNNs), which is jointly trained with a VAE-based architecture, where only relationships prioritised by the learned graph influence feature generation. Due to its relational learning graph-based component, it might be difficult to scale GOGGLE to datasets of higher dimensionality.
Diffusion-based models. TDGGD: The first framework encoding background knowledge about feature relations into diffusion-based models is TDGGD [?]. The relations can express lower and upper bounds over single features or their sum. Rather than explicitly modelling the relations, TDGGD models whether the features are part of such relations or not, which only gives an indication of hidden relations between the columns to the model.
Neural network layer-based models. C-DGM: The first method to constrain deep generative models and guarantee the constraints satisfaction was proposed in [?]. The constraints can capture any set of linear inequalities over the feature space. Differently from the methods surveyed above, this approach relies on a layer to be added right before the sample output layer, which restricts the output space of the model to coincide with the space defined by the constraints. The layer can be added on top of any deep generative model (DGM), and the resulting models are called C-DGMs. For a visual representation on how the layer-based methods can be added on a GAN model, see the right-most schema in Figure 1. As the layer is differentiable and acyclic, it can be added both at inference and training time, while also improving the utility of the model. DGM+DRL: The above layer was further extended in  [?], in order to capture constraints as expressive as disjunctions over linear inequalities (which allow for expressing relations like “if the sum of two features is greater than 10 then the difference of other two features should be lower than 5”), thus modelling non-convex and even disconnected output spaces. The new layer is called Disjunctive Refinement Layer (DRL), and a DGM constructed using DRL is called DGM+DRL. Just as C-DGM, DGM+DRL is able to guarantee alignment with the background knowledge, while also improving utility across all considered baselines.
5 Fidelity
While statistical fidelity is not a good predictor of how well the synthetic data would perform in downstream tasks [?], it can provide useful observations on how the synthetic data compares to the real data. Fidelity was indeed one of the requirements often measured in early works on synthesising tabular data.
5.1 Evaluation Protocol and Metrics
Feature-wise evaluation: compares synthetic and real data feature by feature, using different metrics based on feature type. For continuous features, Wasserstein distance and the Kolmogorov-Smirnov test are common, while Jensen-Shannon divergence or total variation distance are common metrics for discrete features. The difference in treatment is supported empirically in [?], which found that Jensen-Shannon divergence is unstable for measuring the statistical similarity between continuous features’ distributions, especially with no overlap between synthetic and real data.
Pair-wise evaluation: investigates how well relationships between pairs of features are preserved in the synthetic w.r.t. the real data. Common metrics include: (i) RMSE (and MAE) differences in mutual information between pairs of features, and (ii) correlation difference, using Pearson’s coefficient for continuous pairs, Theil’s uncertainty coefficient or the total variation distance for discrete pairs, and the correlation ratio between discrete-continuous pairs of features.
Joint evaluation: compares the joint distribution of the synthetic samples against the real samples’ distribution. It is more difficult to determine how similar are two (potentially high-dimensional) joint distributions. Thus, such an analysis is often conducted in a controlled environment, e.g., on simulated data, rather than real-world data. Relevant evaluation methods include measuring the likelihood fitness score, separately for real and synthetic data (as in [?]).
5.2 Methods
@@
GAN-based models. TGAN [?] is an early GAN-based model that proposes using long short-term memory networks with an attention mechanism to synthesise data sequentially, feature by feature, in order to model tabular data of mixed types. Comparing with conventional statistical models and showing better fidelity performance, TGAN contributed to the popularity of the GAN-based approaches for tabular data. CTGAN: focused on better modelling the types of the variables in an effort to increase fidelity. Indeed, the authors propose a new preprocessing method, which is column-type-specific and models discrete features in the continuous space via Gumbel-Softmax transformations, while it employs mode-specific normalisation via a variational Gaussian mixture model applied for each continuous features. Thanks to these transformations, CTGAN is able to avoid mode collapse, which often burdens other GAN-based models, like medGAN [?] and TableGAN [?]. CTAB-GAN: Building upon CTGAN, CTAB-GAN [?] enhances statistical fidelity by addressing data imbalance and long-tail issues. It introduces a conditional vector to handle mixed-type features and multi-modal continuous variables, extending support beyond discrete and continuous features. Further, it supports mixed-type features: which are features that might have highly-recurring discrete values, but also continuous values across the data points (e.g., a feature Credit Card Transaction USD might often have value 00, but it can also exceed $500$). The benefit of providing support for mixed-type features is reflected in the fidelity performance, with CTAB-GAN outperforming CTGAN. CTAB-GAN+ [?] is an extension of CTAB-GAN, which addresses privacy concerns by training a discriminator using differential private-SGD [?]. Importantly, CTAB-GAN+ demonstrates high fidelity performance.
VAE-based models. TVAE [?], proposed along with CTGAN, was designed to check whether CTGAN’s generator’s lack of access to the real data during training makes it weaker compared to models that do use these data for training their generator, such as models based on Variational Autoencoders (VAEs) [?]. Indeed, TVAE’s architecture showed an improvement over CTGAN in terms of fidelity. However, like CTAB-GAN+, the authors also considered the impact of their proposed models w.r.t. privacy and noted that TVAE’s access to the training data might make it unsuitable in downstream tasks where sensitive data are present.
Normalising flows-based models. Neural Spline Flows: Rather than using the common transformations found in normalising flows, ? [?] proposed using a fully-differentiable module based on monotonic rational-quadratic piecewise functions. This adaptation, combined with the inherent properties of normalising flows, which model the data as the output of an invertible differentiable transformation of a noisy sample drawn from a known distribution, can help synthesise high-fidelity tabular data. The authors note that such a result is conditional on having enough training data compared to the datasets dimensionality.
Diffusion-based models. FinDiff: Motivated by the need to preserve the statistical properties of real data of mixed types in real-world financial contexts, ? [?] proposes FinDiff, which is capable of generating mixed-type tabular financial data using a diffusion model to capture high-dimensional dependencies. Unlike another popular diffusion model, TabDDPM [?], which uses one-hot encoding for categorical features, FinDiff employs an embedding encoding for better handling of mixed data types. This leads to higher fidelity performance compared to the baselines, indicating that FinDiff effectively captures both feature-wise and joint distributions of the data.
Diffusion and AE-based models. AutoDiff: Proposed in [?], AutoDiff is a diffusion-based model designed to preserve statistical fidelity by modelling the joint distribution via an autoencoder, rather than modelling the distributions of individual features separately, as seen in GAN- and diffusion-based models like CTGAN and TabDDPM. More precisely, AutoDiff (i) uses the autoencoder to learn the distribution of the features (which can be of mixed types) in a continuous space, and then (ii) passes these continuous representations to the diffusion model, which generates latent representations that are then decoded back into representations in the original feature space. Moreover, like CTAB-GAN, it introduces an approach to handle mixed-type features by adding a new feature to the autoencoder that encodes the frequency of recurring values in a mixed-type feature.
Diffusion and VAE-based models. TabSyn: Introduced by ? [?], TabSyn trains a score-based diffusion model, like STaSy, but in a joint VAE-learned space of numerical and categorical features. Notably, TabSyn significantly reduces the runtime of diffusion-based models while outperforming baselines w.r.t. fidelity and utility.
6 Privacy
Synthetic data are particularly valuable in privacy-sensitive domains like finance and healthcare. So, there has been growing focus on generating data without revealing sensitive information that could identify entities from the original dataset.
6.1 Evaluation Protocol and Metrics
Privacy is typically evaluated by assessing the risk of attacks succeeding in targeting sensitive information. The most common attacks are: (i) membership attacks, which use ML classifiers or black-box attacks (e.g., see [?]) to determine whether an entity was part of the set used to train a model, (ii) attribute disclosure attacks, which try to gain access to sensitive attributes of an entity from the real dataset, and (iii) re-identification attacks, which try to map a synthetic data point back to the original dataset. To measure the risk of the attacks being successful, for (i), AUCROC is reported for the trained attack models (typically one for each class of a feature of interest), for (ii), precision and sensitivity (i.e., recall) are measured while varying the number of attributes known to the attacker, and, for (iii), common metrics include: the distance to the closest record (DCR) [?] (computed as the minimum L2 distance from a synthetic data point to each real data point), the nearest neighbour distance ratio (computed as the ratio between the distances to the closest and second-closest real neighbours for a synthetic record), but also k-anonymisation, delta presence, and the identifiability score, as seen in [?].
Finally, methods using differential privacy [?] introduce noise during training to reduce retention of original data features. Privacy is then evaluated using different privacy budget values, which determine the amount of added noise and balance utility and privacy (e.g., see [?]).
6.2 Methods
AE and GAN-based models. medGAN: Proposed by ? [?] for synthesising healthcare datasets, medGAN is one of the pioneering models for synthesising tabular data using deep generative models. medGAN generates distributed representations of health records using the generator of a GAN-based model, then decodes these representations using an autoencoder pretrained on the real data, and passes the decoded representations to the discriminator component of the GAN, which is trained to distinguish the synthetic from the real data. In terms of privacy, medGAN showed low attribute disclosure risks, except when an attacker focused on a small number of data points. IT-GAN: Similar to OCT-GAN but designed to enhance privacy robustness, IT-GAN [?] employs a generator based on neural ODEs to balance utility and privacy protection. Instead of directly synthesising samples in the input space, the generator produces hidden representations, as seen also in medGAN. Notably, it ensures that the hidden representation space matches the input dimensionality, enabling the application of ODEs.

### v2

@@ 2 Synthesising Tabular Data
5. Diversity requirement. Synthetic data should cover the full variability of the real data distribution. While closely related to fidelity, diversity focuses on a different aspect of the data distribution. Specifically, diversity ensures that the model does not omit regions of the output sample space, which is what would happen in the case of a mode collapse, for example. Continuing the earlier example, it is likely that the WiDS dataset contains under-represented patient groups, such as young adults with specific co-morbidities or rare physiological responses, e.g., captured as extreme but valid laboratory test result values. In such cases, failing to meet the diversity requirement might only produce “average” patient profiles. Thus, the goal is to learn parameters $\theta$ such that the support of $p_{\theta}$ fully covers the support of $p_{X}$, which would allow for rare feature values or combinations of values to be synthesised (with appropriate probability).
Table 2: Overview of the evaluation metrics typically used for the five synthetic tabular data requirements. The formula and terminology definitions for each metric are provided in Appendix A.
Requirement Evaluation Metric
Utility
Accuracy
F1-score
Root mean square error (RMSE)
Explained variance
Epistemic parity
Alignment
Constraint violation rate (CVR)
Constraint violation coverage (CVC)
Sample-wise constraint violation coverage (sCVC)
Fidelity Feature-wise
Wasserstein distance
Kolmogorov-Smirnov test statistic
Jensen-Shannon divergence
Total variation distance
Pair-wise
RMSE differences in mutual information
MAE differences in mutual information
Correlation difference
Joint
Likelihood fitness score
$\alpha$-Precision
Density
Clipped density
Privacy
AUCROC (for membership attacks)
Precision (for attribute disclosure attacks)
Recall/Sensitivity (for attribute disclosure attacks)
Distance to the closest record (DCR)
Nearest neighbour distance ratio (NNDR)
k-anonymity
Diversity
$\beta$-Recall
Coverage
Clipped coverage
3 Requirements and Evaluation Metrics
To systematically evaluate the quality of synthetic tabular data and guide the selection of appropriate synthesisers, in this section, we detail each of the five requirements, along with the specific evaluation protocols and metrics used to quantify them, and discuss their significance in real-world applications.
For each of the five requirements above, Table 2 lists the metrics that are commonly used to evaluate the generated data. Further, in Appendix A, we provide the formula for each of these metrics, along with the terminology definitions.
3.1 Utility
Determining whether synthetic data can replace real training data in downstream tasks has been the most common evaluation approach (e.g., [Choi et al., 2017; Park et al., 2018; Kotelnikov et al., 2023; Borisov et al., 2023]). The main reason for this is that synthetic data are often generated to augment or create datasets for real-world applications, including high-stakes fields like finance and healthcare. In these scenarios, generating synthetic samples for rare events, e.g., fraudulent transactions and specific medical diagnoses [Matthys et al., 2021], can help with class balancing and ultimately improve the quality of the predictions in downstream tasks. Therefore, utility still remains a key requirement for synthetic data and benchmark for comparing synthesisers.
Evaluation Protocol and Metrics.
The most common protocol used to evaluate the utility of generative models is the Train on Synthetic, Test on Real (TSTR) Protocol (see, e.g., [Esteban et al., 2017]). The procedure is to train a suite of ML models (e.g., a support vector machine, XGBoost) on the synthetic data, and then test their performance on a real test set, reporting the performance using different metrics on the ground of the task defined by the dataset of interest: for classification (resp., regression) datasets, standard metrics include accuracy and F1 score (resp., root mean square error and explained variance).
Beyond predictive performance (TSTR), foundational work on classical statistical, non-deep learning methods for generating fully synthetic data in [Raghunathan et al., 2003] aimed to preserve the users’ ability to obtain valid statistical inferences without compromising the confidentiality of survey respondents. Further, in the specific case of differentially private settings, recent work by Rosenblatt et al. [2023] similarly argues that synthetic data must also support valid inference, ensuring that statistical conclusions derived from real data are preserved. To measure this, they introduce a metric called epistemic parity, which evaluates whether empirical conclusions of peer-reviewed papers on real, publicly available data can be reproduced on synthetic datasets. To this end, in the context of synthesising data intended for scientific discovery, Montoya Perez et al. [2024] emphasise that evaluating the validity and power of statistical tests is essential.
3.2 Alignment
Brought to the attention of the research community more recently [Chen et al., 2019] than the other requirements, background knowledge alignment is an important condition for synthetic data when evaluating its realism. This requirement is important in practice, as many real-world applications have domain-specific knowledge that the synthetic data must satisfy. For example, synthetic patient data in healthcare must adhere to physiological constraints, ensuring a recorded minimum value for an indicator (like blood pressure or haemoglobin level) is not higher than the maximum value recorded. Similarly, in finance, it is important for generated market data to comply with known economic constraints. In resource management, synthetic inventory and demand data must not violate logistical constraints, such as production capacity limits or non-negative stock levels.
Evaluation Protocol and Metrics.
In [Stoian et al., 2024], three metrics have been proposed to evaluate alignment: (i) the Constraint Violation Rate (CVR), which computes the percentage of samples that do not satisfy at least one of the constraints in the available set of constraints, (ii) Constraint Violation Coverage (CVC), computing the percentage of constraints that have been violated at least once by the sample set, and (iii) the sample-wise Constraint Violation Coverage (sCVC) which determines the average percentage of samples violating each of the constraints. Out of these three metrics, CVR has also been used in [Jia et al., 2024], albeit called feasibility rate there.
3.3 Fidelity
While statistical fidelity is not a good predictor of downstream task performance [Hansen et al., 2023], it can provide useful observations on how the synthetic data compares to the real data. Fidelity is indeed one of the requirements often evaluated in early works on synthesising tabular data and it is crucial for applications such as economic modelling and census data release. In these areas, preserving the statistical properties of the real data distribution, like demographics, allows for accurate downstream sociological and economic analyses. Similarly, generating data that captures the statistical complexity of real-world data is valuable in other domains, such as the robust testing of data processing software.
Evaluation Protocol and Metrics.
  1. 1.
    Feature-wise evaluation: compares synthetic and real data feature by feature, using different metrics based on feature type. For continuous features, Wasserstein distance and the Kolmogorov-Smirnov test statistic are common, while Jensen-Shannon divergence or total variation distance are common metrics for discrete features. The difference in treatment is supported empirically in [Zhao et al., 2021], which found that Jensen-Shannon divergence is unstable for measuring the statistical similarity between continuous features’ distributions, especially with no overlap between synthetic and real data.
  2. 2.
    Pair-wise evaluation: investigates how well relationships between pairs of features are preserved in the synthetic w.r.t. the real data. Common metrics include: (i) RMSE (and MAE) differences in mutual information between pairs of features, and (ii) correlation difference, using Pearson’s coefficient for continuous pairs, Theil’s uncertainty coefficient for discrete pairs, and the correlation ratio between discrete-continuous pairs of features.
  3. 3.
    Joint evaluation: compares the joint distribution of the synthetic samples against the real samples’ distribution. It is more difficult to determine how similar are two (potentially high-dimensional) joint distributions. Thus, such an analysis is often conducted in a controlled environment, e.g., on simulated data, rather than real-world data. Relevant evaluation methods include measuring: (i) the likelihood fitness score, separately for real and synthetic data (as in [Xu et al., 2019]), (ii) $\alpha$-precision [Alaa et al., 2022], which is a variant of precision representing the fraction of synthetic samples that resemble the most typical fraction $\alpha$ of real samples, (iii) density (defined in [Naeem et al., 2020]), which improves upon the standard precision metric by counting the number of real manifolds each synthetic sample falls within, and (iv) clipped density, which has been proposed by Salvy et al. [2025] as an improvement on the density metric in terms of robustness, clipping the radius of the nearest-neighbour manifolds used to measure the density to the median of the distances to the $k$-th nearest neighbour.
3.4 Privacy
Synthetic data are particularly valuable in privacy-sensitive domains. Thus, there has been growing focus on generating data without revealing sensitive information that could identify entities from the original dataset. For example, sharing synthetic patient records for medical research allows for collaboration without risking the re-identification of individuals or the disclosure of sensitive health information. Similarly, privacy is an important requirement in education analytics, where synthetic educational data allows for studying student performance trends without exposing individual student identities or sensitive background information. Finance is another key domain where privacy is a concern and synthesising transaction or credit data to be released for academic or commercial research must be done in a way that protects customer confidentiality.
Evaluation Protocol and Metrics.
Privacy is typically evaluated by assessing the risk of attacks succeeding in targeting sensitive information. The most common attacks are: (i) membership inference attacks, which use ML classifiers or black-box attacks (e.g., see [Chen et al., 2020]) to determine whether an entity was part of the set used to train a model, (ii) attribute disclosure attacks, which try to gain access to sensitive attributes of an entity from the real dataset, and (iii) re-identification attacks, which try to map a synthetic data point back to the original dataset. To measure the risk of the attacks being successful, for (i), AUCROC is often reported for the trained attack models (typically one for each class of a feature of interest). However, this metric can be misleading, as it might suggest that a model is secure even if the privacy of a few users has been confidently breached or, conversely, imply that an attack is successful based on unreliable scores. To this end, Carlini et al. [2022] argue that membership inference attacks should be evaluated using the true-positive rate (TPR) at low false-positive rates (FPR), which is capable of detecting such high-confidence attacks that aggregated metrics, such as AUCROC, might overlook. For (ii), precision and sensitivity (i.e., recall) are measured while varying the number of attributes known to the attacker, and, for (iii), common metrics include: the distance to the closest record (DCR) [Zhao et al., 2021] (computed as the minimum L2 distance from a synthetic data point to each real data point), the nearest neighbour distance ratio (computed as the ratio between the distances to the closest and second-closest real neighbours for a synthetic record), but also k-anonymisation, delta presence, and the identifiability score, as seen in [Jia et al., 2024].
Finally, methods using differential privacy [Dwork, 2006] introduce noise during training to reduce retention of original data features. Privacy is then evaluated using different privacy budget values, which determine the amount of added noise and balance utility and privacy (e.g., see [Park et al., 2018]).
3.5 Diversity
While fidelity ensures that generated samples resemble the real ones, diversity ensures that the synthetic data cover the full variability of the real data distribution [Naeem et al., 2020]. This plays an important role in avoiding mode collapse, where a synthesiser produces samples that fit well within the data distribution but which are repetitive, failing to represent rare feature combinations or outliers which are present in the real data. This is of particular importance in healthcare as described earlier, but also in domains such as finance, e.g., for detecting fraud patterns.
Evaluation Protocol and Metrics.
Typically, diversity is evaluated by determining the proportion of the real data manifold that is represented by the generated data. There are multiple ways to compute this, with the most common metrics being: (i) $\beta$-recall [Alaa et al., 2022], which is a variant of recall evaluating whether the synthetic data covers the variability of the real data (e.g., as used in [Liu et al., 2022]), (ii) coverage (defined in [Naeem et al., 2020] and used in [Kim et al., 2023; Kotelnikov et al., 2023]), which improves upon the standard recall metric by building nearest neighbour manifolds around the real samples, instead of around the synthetic samples, to ultimately measure the proportion of real samples whose neighbourhoods contain at least one synthetic sample, and (iii) clipped coverage, which has been proposed by Salvy et al. [2025] as an improvement on the coverage metric in terms of robustness (e.g., unlike the standard metric which only checks for the presence of synthetic samples within the real manifolds, clipped coverage counts them and bounds their contributions to effectively reflect the real data distribution rather than just its support).
3.6 Relationships among the Requirements
As emphasised by Naeem et al. [2020], assessing a generative model is an inherently difficult task, as no single metric can capture all the desired properties of synthetic data. And, while the requirements described in Sections 3.1-3.5 are presented individually, they are not orthogonal. In fact, these requirements’ relationships vary from mutual reinforcement to mutual tension, possibly leading to tradeoffs. Thus, forming an understanding of these relationships is crucial for practitioners to select the appropriate synthesiser for their specific use case.
Fidelity and Utility.
Intuitively, higher statistical fidelity, where the distribution of the synthetic data $p_{\theta}$ closely approximates the distribution of the real data $p_{X}$, often yields higher utility. However, this is not always the case, as a high statistical fidelity does not guarantee optimal utility in downstream tasks and vice-versa. Indeed, as shown by Hansen et al. [2023], fidelity alone is insufficient for assessing the synthetic data’s utility, and should not be used as an indicator for it.
Alignment and Fidelity.
The relationship between these two requirements can be seen as hierarchical. A model with “perfect” fidelity would implicitly satisfy all background knowledge constraints present in the real data, assuming such data do not contain any errors. However, high alignment does not necessarily imply high fidelity. In spite of this, alignment plays a vital role for any model that is to be deployed in domains where constraints must be guaranteed [Stoian et al., 2024], whereas fidelity often remains an approximation.
Furthermore, it is still possible to have real data that contains erroneous values (e.g., swapped minimum and maximum values due to human error). In such cases, the constraints can be used to validate and correct the data before synthesis. Indeed, if these errors are not addressed, a model that strictly satisfies the alignment requirement may inevitably deviate from statistical fidelity measured in this case against the corrupted real data.
The Privacy Trade-off (Privacy, Fidelity, and Utility).
The tension between privacy and utility is perhaps the most critical trade-off in tabular data synthesising. In particular, any method that claims to be preserving privacy must balance privacy and utility jointly rather than separately Montoya Perez et al. [2024]. To satisfy differential privacy, for example, noise must be added during sample generation. And, as more such noise is added (i.e., to better protect the data subjects), statistical fidelity decreases, which often leads to lower downstream performance (i.e., utility). It is thus important for synthesising methods to determine the acceptable extent of utility to be sacrificed in order to achieve robust privacy guarantees.
Diversity and Fidelity.
Intuitively, fidelity and diversity are associated with precision and recall, respectively [Sajjadi et al., 2018; Naeem et al., 2020]. Models suffering from mode collapse may still produce high-fidelity samples (i.e., high-precision), but fail to cover the full variability of the real data distribution (i.e., low diversity or low recall). Conversely, maximising diversity may result in generating outliers or noisy samples, which often leads to lower fidelity. Thus, these two metrics are seen as complementary trade-offs, as maximising one without the other leads to suboptimal synthetic data. Further, in an effort to disentangling fidelity from diversity, different metrics have been proposed to measure each of the two requirements, which were often improved precision and recall variants as first presented in [Kynkäänniemi et al., 2019]), and subsequently in Naeem et al. [2020]; Alaa et al. [2022]). Nevertheless, the recent position paper by Räisä et al. [2025] argues that current metrics for measuring fidelity and diversity are flawed and advocates for metrics with meaningful absolute values, rather than just relative comparisons, to truly determine if a model is useful in practice.
@@ 3.6 Relationships among the Requirements
4 Methods
Having established the key requirements for synthetic tabular data, we now survey the deep generative models designed to meet them. To provide a clear structural overview, we categorise these approaches based on their underlying generative paradigm: generative adversarial networks (GANs), diffusion models, transformers and large language models (LLMs), variational autoencoders (VAEs), constraints-enforcing methods, along with other deep learning methods (e.g., approaches stemming from other domains but which have been adapted to tabular data synthesis). Finally, we also discuss briefly non-deep learning alternatives.
4.1 Generative Adversarial Network-based Methods
4.1.1 GAN-based models
@@
Many of the works surveyed here are based on the original Generative Adversarial Network (GAN) [Goodfellow et al., 2014], which relies on two neural networks: (i) a generator, which given a noise vector returns a synthetic datapoint and hence learns the real data distribution, and (ii) a discriminator, which given a data point (either real or synthetic) classifies it as either real or synthetic and hence learns to distinguish between synthetic and real data. The GAN jointly trains these two models in an adversarial framework, where the generator’s objective is to produce samples that confuse the discriminator. See the first schema from the left in Figure 2 for a simple abstraction of GAN-based models.
CWGAN: Building on the seminal Wasserstein GAN (WGAN) [Arjovsky et al., 2017], which used the Wasserstein distance as a loss function for a GAN model, Engelmann & Lessmann [2022] proposed CWGAN as an oversampling framework for class balancing. CWGAN generates synthetic samples for underrepresented classes, augmenting the training set. Results indicate that oversampling is beneficial primarily for strongly non-linear datasets. OCT-GAN: proposed by Kim et al. [2021], incorporates neural ordinary differential equations (neural ODEs) in their GAN through the introduction of a new layer that allows for the extraction of a sequence of hidden vector representations given an input sample, i.e., the hidden evolution trajectory of the sample. The trajectory is used to help the discriminator decide whether a sample is real or synthetic. While improving utility, the usage of ODEs makes OCT-GAN slower than most GAN-based models. TGAN [Xu & Veeramachaneni, 2018] is an early GAN-based model that proposes using long short-term memory networks with an attention mechanism to synthesise data sequentially, feature by feature, in order to model tabular data of mixed types. Comparing with conventional statistical models and showing better fidelity performance, TGAN contributed to the popularity of the GAN-based approaches for tabular data.
@@
CTGAN: focused on better modelling the types of the variables in an effort to increase fidelity. Indeed, the authors propose a new preprocessing method, which is column-type-specific and models discrete features in the continuous space via Gumbel-Softmax transformations, while it employs mode-specific normalisation via a variational Gaussian mixture model applied for each continuous features. Thanks to these transformations, CTGAN is able to avoid mode collapse, which often burdens other GAN-based models, like medGAN [Choi et al., 2017] and TableGAN [Park et al., 2018]. CTAB-GAN: Building upon CTGAN, CTAB-GAN [Zhao et al., 2021] enhances statistical fidelity by addressing data imbalance and long-tail issues. It introduces a conditional vector to handle mixed-type features and multi-modal continuous variables, extending support beyond discrete and continuous features. Further, it supports mixed-type features: which are features that might have highly-recurring discrete values, but also continuous values across the data points (e.g., a feature Credit Card Transaction USD might often have value 0, but it can also exceed $500$). The benefit of providing support for mixed-type features is reflected in the fidelity performance, with CTAB-GAN outperforming CTGAN. CTAB-GAN+ [Zhao et al., 2022] is an extension of CTAB-GAN, which addresses privacy concerns by training a discriminator using differential private-SGD [Abadi et al., 2016]. Importantly, CTAB-GAN+ demonstrates high fidelity performance.